OpenStudy (anonymous):

What is the surface area of 0.5 cm 1.0 cm 1.5 cm 2.0 cm Thanks :)

OpenStudy (anonymous):

ques is nt complete

zepdrix (zepdrix):

Is there a picture to follow cent? :D

OpenStudy (anonymous):

No its for a school project and you had to cut an egg into four cubes and the measurements of the cubes are supposed to be those above.

zepdrix (zepdrix):

So you cut an egg into 4 cubes, and widdled the smallest cube down to a size of .5 on each side. And did the same for the other measurements i guess..? :X And so the problem is, find the surface area of each cube with the given side length? :D

OpenStudy (anonymous):

Yeah sorry I forgot to put that :) And then it has this its a chart and this is the part I need to fill out. Surface Area (cm2) (Side x Side x 6)

zepdrix (zepdrix):

Hmm well if it's a cube, then each of it's sides are suppose to have the SAME measure as one another. So if one side is 0.5cm, then the length, and width of that particular cube will be 0.5cm also! :D \[\large (0.5cm \times 0.5cm \times 6) = 1.5cm^2\] So this is what the first one would look like, understand? :D

OpenStudy (anonymous):

Yup thanks so much. For the next one it would be (1.0cm X 1.0cm X 6) = 6?

zepdrix (zepdrix):

Yah looks good :)

zepdrix (zepdrix):

You're multiplying centimeters and centimeters, so you should end up with square centimeters for your units, not a big deal, but your teacher might want to see the units ^^ (Unless it's labeled on your chart already) :D

OpenStudy (anonymous):

Fabulous thanks so much:)

OpenStudy (anonymous):

Ah im sorry could you help me with one more?

zepdrix (zepdrix):

sup :D

OpenStudy (anonymous):

Alright so this is the next part of the chart. We need to find the Volume (cm3) (side x side x side). Would we just do 0.5 X 3 = 1.5cm^3?

zepdrix (zepdrix):

No not quite, \[(0.5cm \times 0.5cm \times 0.5cm) \neq (0.5cm \times 3)\] Don't use the 3 :D Use the equation on the left.

OpenStudy (anonymous):

0.375cm?

OpenStudy (anonymous):

Lol sorry im confused

OpenStudy (anonymous):

0.125cm3

zepdrix (zepdrix):

If (side)=0.5cm then,\[\large (side \times side \times side)=(0.5cm \times 0.5cm \times 0.5cm)=?\]

zepdrix (zepdrix):

Yah i think thats right c: gj

OpenStudy (anonymous):

Thanks :) And then I have one more part of the chart. Are you up for it?

zepdrix (zepdrix):

LOL XD I dunno, just post it :3 i'll see what i can do

OpenStudy (anonymous):

Alright last one thanks soo much. Surface Area to Volume Ratio (surface area/vol.)

OpenStudy (anonymous):

Do I just divide Surface Area to the Volume Ratio?

zepdrix (zepdrix):

Hmmmm yah that sounds right :D What'd you get for the 0.5cm cube?

OpenStudy (anonymous):

When I divided 0.5/0.125 I got 12 then do I divide the the 2cm and the 3cm?

zepdrix (zepdrix):

You divided 1.5/0.125, and got 12 :D (surface area/volume) Hmm yah that looks right. I guess that tells us the ratio of the surface area to the volume. The tiny 0.5cm cube has 12 times as much surface area as it does volume.

OpenStudy (anonymous):

Do I just put 12 or do I divide the cm part like the 2cm and the 3cm?

zepdrix (zepdrix):

Hmmmm good question. \[(\frac{ 1.5cm^2 }{ 0.125cm^3 })=\frac{ 12 }{ cm }\] Hmm that's kind of awkward looking :O The surface area is 12 units larger than the volume per centimeter, .. i guess? :3 hmm sorry i dunno on that part :D mind blanking.

OpenStudy (anonymous):

Aha it's alright. I'll just go with putting 12. Thanks so much :o)